Lionblaze15:

HELPPPPPPP A flowerpot with a mass of 3 kg is sitting on a windowsill 30 m above the ground. How much potential energy does the flowerpot have? Group of answer choices 882 J 2700 J 270 J 90 J

DCharger96:

I would suggest that it is 90 J

Lionblaze15:

That was wrong last time :(

Lionblaze15:

Seth do u know?

sethmorrison03:

270j?

Lionblaze15:

Thxs I'll try it

justjm:

P = mgh m=3 kg g = acceleration d/t gravity = 9.81 m/s^2 h= 30 m P= (3 kg)(9.81 m/s^2)(30 m) = approx 882.9 J A is the closest answer